Sandra is 36 years old. her son is 14 years old. how many years ago was Sandra 3 times older than her son?

Found 2 solutions by jim_thompson5910, stanbon:
Answer by jim_thompson5910(35256)

Let x = the number of years


If we subtract x from each age, we get


Sandra's age x yrs ago: 36 - x
Son's age x yrs ago: 14 - x


We want to know when she was 3 times older, so


(Sandra's age x years ago) = 3*(Son's age x yrs ago)


36 - x = 3*(14 - x)


Now we solve for x


36 - x = 3*(14 - x)


36 - x = 3*(14) + 3*(-x)


36 - x = 42 - 3x


36 = 42 - 3x + x


36 - 42 = -3x + x


-6 = -2x


-6/(-2) = x


3 = x


x = 3


So 3 years ago, Sandra was 3 times older than her son (sandra = 33 years, son = 11 years)

Answer by stanbon(75887)

Sandra is 36 years old. her son is 14 years old. how many years ago was Sandra 3 times older than her son?
------
Let the number of years be "x"::
Equation:
36-x = 3(14-x)
------
36-x = 42-3x
-x + 3x = 42-36
2x = 6
x = 3 years ago
-----------------------
Checking::
36-3 = 3(14-3)
33 = 3(11)
Cheers,
Stan H.
--------------